OpenStudy (rockycatsrule):

2. (3 pt) Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 72 when x = 6. A. y = x2 + 25 B. y = 2x2 C. y = 4x2 D. y = 12x

OpenStudy (welshfella):

Direct variation is y = k x^2 where k is some constant Plug in y = 72 and x = 6 into the above and solve for k Then the requred equation is y = k x^2 (insert the found value of k)

OpenStudy (rockycatsrule):

thx for helping me

OpenStudy (tkhunny):

A. - Constant Term - BAD y = x2 + 25 B. y = 2x2 C. y = 4x2 D. - No Squared Term - BAD y = 12x Looks like it's down to B or C.

OpenStudy (anonymous):

1. Any equation that does not have an exponential value for X is a linear one. Therefore only option D is non linear 2. Y is 3 times the value of X; Therefore Y=3X 3. A number times 3 squared =36; A number times3x3=36, A number=36/9, A number = 4. Therefore Y=4X squared. 4> Inverse: As one increases, the other value decreases. Option A with X values of 2345 and Corresponding Y value of Y 30, 20, 15, 12,. We ask a question what value can we divide 2 into to obtain Y value of 30 : What value/2=30; what value =2 x 30; what value =60. Therefore f(X0 =60/X; We test, X value of 3; 60/3=20; X value of 4; 60/4=15; X value of 5; 60/5=12 We vsee only option A will fit this pattern. 5. When X =2, Y will equal 5 ; Therefore we can say : What value /X=5; Replacing X with 2, we can say what value/2=5; Multiplying both side by 2 we have 2x what value,2=2x5 ; 2/2=1 so we can say 1 Times what value =10, We can say The value =10 ; Therefore we can say: 10(what value)/2(X)=5 and 10/X=Y or Y=10/X I've gone about this in a slow methodically way to show the reasoning to help you remember. Hope this helps. Good Luck